\newcommand{\etalchar}[1]{$^{#1}$}
\begin{thebibliography}{FBH{\etalchar{+}}22}

\bibitem[AIH24]{AIH24}
Yamato Arai, Yuma Ichikawa, and Koji Hukushima.
\newblock Adaptive flip graph algorithm for matrix multiplication.
\newblock In {\em Proceedings of the 2024 International Symposium on Symbolic
  and Algebraic Computation}, ISSAC '24, page 292–298, New York, NY, USA,
  2024. Association for Computing Machinery.

\bibitem[Ale97]{Ale97}
V.~B. Alekseyev.
\newblock Maximal extensions with simple multiplication for the algebra of
  matrices of the second order.
\newblock {\em Discrete Mathematics and Applications}, 7(1):89--102, 1997.

\bibitem[BC85]{BC85}
Werner Büchi and Michael Clausen.
\newblock On a class of primary algebras of minimal rank.
\newblock {\em Linear Algebra and its Applications}, 69:249--268, 1985.

\bibitem[BCS97]{BCS97}
Peter B{\"u}rgisser, Michael Clausen, and M.~Amin Shokrollahi.
\newblock {\em Algebraic complexity theory}, volume 315.
\newblock Springer Science \& Business Media, 1997.

\bibitem[BD73]{BD73}
Roger~W. Brockett and David Dobkin.
\newblock On the optimal evaluation of a set of bilinear forms.
\newblock In {\em Proceedings STOC}, STOC '73, page 88–95, 1973.

\bibitem[BILR19]{BILR19}
Grey Ballard, Christian Ikenmeyer, Joseph~M Landsberg, and Nick Ryder.
\newblock The geometry of rank decompositions of matrix multiplication {I}{I}:
  $3 \times 3$ matrices.
\newblock {\em Journal of Pure and Applied Algebra}, 223(8):3205--3224, 2019.

\bibitem[Bre70]{Bre70}
Richard~P Brent.
\newblock Algorithms for matrix multiplication.
\newblock Tech. Report STAN-CS-70-157, Stanford University, 1970.

\bibitem[Cha86]{Cha86}
Philippe Chatelin.
\newblock On transformations of algorithms to multiply $2 \times 2$ matrices.
\newblock {\em Information Processing Letters}, 22(1):1--5, 1986.

\bibitem[CILO19]{CILO19}
Luca Chiantini, Christian Ikenmeyer, J.~M. Landsberg, and Giorgio Ottaviani.
\newblock The geometry of rank decompositions of matrix multiplication {I}: $2
  \times 2$ matrices.
\newblock {\em Experimental Mathematics}, 28(3):322--327, 2019.

\bibitem[Cla88]{Cla88}
M~Clausen.
\newblock Beitr\"age zum {E}ntwurf schneller {S}pektraltransformationen.
\newblock Habilitation Universit\"at Karlsruhe, 1988.

\bibitem[dG78]{deG78}
Hans~F. de~Groote.
\newblock On varieties of optimal algorithms for the computation of bilinear
  mappings {I}{I}. optimal algorithms for $2 \times 2$-matrix multiplication.
\newblock {\em Theoretical Computer Science}, 7(2):127--148, 1978.

\bibitem[DIS11]{DIS11}
Charles~\'{E}ric Drevet, Md.~Nazrul Islam, and \'{E}ric Schost.
\newblock Optimization techniques for small matrix multiplication.
\newblock {\em Theoretical Computer Science}, 412(22):2219--2236, 2011.

\bibitem[FBH{\etalchar{+}}22]{Fawetal22}
Alhussein Fawzi, Matej Balog, Aja Huang, Thomas Hubert, Bernardino
  Romera-Paredes, Mohammadamin Barekatain, Alexander Novikov, Francisco J.~R.
  Ruiz, Julian Schrittwieser, Grzegorz Swirszcz, David Silver, Demis Hassabis,
  and Pushmeet Kohli.
\newblock Discovering faster matrix multiplication algorithms with
  reinforcement learning.
\newblock {\em Nature}, 610(7930):47--53, 2022.

\bibitem[Fid72]{Fid72}
Charles~M. Fiduccia.
\newblock On obtaining upper bounds on the complexity of matrix multiplication.
\newblock In {\em Proceedings Complexity of Computer Computations}, pages
  31--40. Springer, 1972.

\bibitem[Gas71]{Gas71}
No{\"e}l Gastinel.
\newblock Sur le calcul des produits de matrices.
\newblock {\em Numerische Mathematik}, 17(3):222--229, 1971.

\bibitem[GK00]{GK00}
Ann~Q Gates and Vladik Kreinovich.
\newblock Strassen's algorithm made (some-what) more natural: A pedagogical
  remark.
\newblock Technical Report 502, University of Texas at El Paso, Department of
  Computer Science, 2000.

\bibitem[GM17]{GM17}
Joshua~A Grochow and Cristopher Moore.
\newblock Designing {S}trassen's algorithm.
\newblock {\em arXiv:1708.09398}, 2017.

\bibitem[HK71]{HK71}
John~E Hopcroft and Leslie~R Kerr.
\newblock On minimizing the number of multiplications necessary for matrix
  multiplication.
\newblock {\em SIAM Journal on Applied Mathematics}, 20(1):30--36, 1971.

\bibitem[HKS19]{HKS19}
Marijn~JH Heule, Manuel Kauers, and Martina Seidl.
\newblock Local search for fast matrix multiplication.
\newblock In {\em Theory and Applications of Satisfiability Testing--SAT 2019},
  pages 155--163. Springer, 2019.

\bibitem[HKS21]{HKS21}
Marijn~JH Heule, Manuel Kauers, and Martina Seidl.
\newblock New ways to multiply $3 \times 3$-matrices.
\newblock {\em Journal of Symbolic Computation}, 104:899--916, 2021.

\bibitem[IL19]{IL19}
C.~Ikenmeyer and V.~Lysikov.
\newblock Strassen's $2 \times 2$ matrix multiplication algorithm: a conceptual
  perspective.
\newblock {\em Annali dell'Universit{\`a} di Ferrara}, 65:241--248, 2019.

\bibitem[JM86]{JM86}
Rodney~W. Johnson and Aileen~M. McLoughlin.
\newblock Noncommutative bilinear algorithms for {$3\times 3$} matrix
  multiplication.
\newblock {\em SIAM J. Comput.}, 15(2):595--603, 1986.

\bibitem[KM23]{KM23}
Manuel Kauers and Jakob Moosbauer.
\newblock Flip graphs for matrix multiplication.
\newblock In {\em Proceedings of the 2023 International Symposium on Symbolic
  and Algebraic Computation}, pages 381--388, 2023.

\bibitem[KM25]{KM25}
Manuel Kauers and Jakob Moosbauer.
\newblock Some new non-commutative matrix multiplication algorithms of size
  ($n$, $m$, 6).
\newblock {\em ACM Commun. Comput. Algebra}, 58:1–11, January 2025.

\bibitem[Lad76]{Lad76}
Julian~D Laderman.
\newblock A noncommutative algorithm for multiplying $3 \times 3$ matrices
  using 23 multiplications.
\newblock {\em Bull. Amer. Math. Soc.}, 82(1):126--128, 1976.

\bibitem[Laf75]{Laf75}
Jean-Claude Lafon.
\newblock Optimum computation of $p$ bilinear forms.
\newblock {\em Linear Algebra and its Applications}, 10(3):225--240, 1975.

\bibitem[Lan08]{Lan08}
J~Landsberg.
\newblock Geometry and the complexity of matrix multiplication.
\newblock {\em Bulletin of the American Mathematical Society}, 45(2):247--284,
  2008.

\bibitem[Lic85]{Lic85}
Thomas Lickteig.
\newblock Typical tensorial rank.
\newblock {\em Linear algebra and its applications}, 69:95--120, 1985.

\bibitem[Mak75]{Mak75}
O.M. Makarov.
\newblock The connection between two multiplication algorithms.
\newblock {\em USSR Computational Mathematics and Mathematical Physics},
  15(1):218--223, 1975.

\bibitem[Mak87]{Mak87}
O.M. Makarov.
\newblock A non-commutative algorithm for multiplying 5 × 5 matrices using one
  hundred multiplications.
\newblock {\em USSR Computational Mathematics and Mathematical Physics},
  27(1):205--207, 1987.

\bibitem[MP25]{MP25}
Jakob Moosbauer and Michael Poole.
\newblock Flip graphs with symmetry and new matrix multiplication schemes.
\newblock arXiv:2502.04514, 2025.

\bibitem[Pat74]{Pat74}
Mike Paterson.
\newblock Complexity of product and closure algorithms for matrices.
\newblock In {\em Proceedings of the International Congress of Mathematicians},
  volume~2, pages 483--489, 1974.

\bibitem[Pat09]{Pat09}
Mike Paterson.
\newblock Strassen's symmetries.
\newblock Presentation at Leslie Valiant's 60th birthday celebration, May 2009.

\bibitem[Sed17]{Sed17}
Alexandre Sedoglavic.
\newblock A non-commutative algorithm for multiplying ($7 \times 7$) matrices
  using 250 multiplications.
\newblock arXiv:1712.07935, 2017.

\bibitem[Sed25]{Sed25}
Alexandre Sedoglavic.
\newblock Yet another catalogue of fast matrix multiplication algorithms.
\newblock \url{https://fmm.univ-lille.fr}, 2025.

\bibitem[Smi13]{Smi13}
Alexey~V Smirnov.
\newblock The bilinear complexity and practical algorithms for matrix
  multiplication.
\newblock {\em Computational Mathematics and Mathematical Physics},
  53:1781--1795, 2013.

\bibitem[Smi21]{Smi21}
Alexey Smirnov.
\newblock Several bilinear algorithms for matrix multiplication problems
  $\left<3, {P}, {Q}\right>$, April 2021.
\newblock
  \url{https://www.researchgate.net/publication/350897049_Several_Bilinear_Algorithms_for_Matrix_Multiplication_Problems_3_P_Q}.

\bibitem[Smi23]{Smi23}
Alexey Smirnov.
\newblock Bilinear algorithm $\left<4, 4, 6; 73\right>$ for matrix
  multiplication, January 2023.
\newblock
  \url{https://www.researchgate.net/publication/366989954_Bilinear_Algorithm_4_4_6_73_for_Matrix_Multiplication}.

\bibitem[SS21]{SS21}
Alexandre Sedoglavic and Alexey~V. Smirnov.
\newblock The tensor rank of $5 \times 5$ matrices multiplication is bounded by
  98 and its border rank by 89.
\newblock In {\em Proceedings of the 2021 International Symposium on Symbolic
  and Algebraic Computation}, ISSAC '21, page 345–351, 2021.

\bibitem[Str69]{Str69}
Volker Strassen.
\newblock Gaussian elimination is not optimal.
\newblock {\em Numerische Mathematik}, 13(4):354--356, 1969.

\bibitem[Yuv78]{Yuv78}
Gideon Yuval.
\newblock A simple proof of {S}trassen's result.
\newblock {\em Information Processing Letters}, 7(6):285--286, 1978.

\end{thebibliography}

\documentclass[a4paper]{article}
\usepackage{graphicx,amsmath,amssymb,amsthm,mathtools,parskip,url,tikz,hyperref}
\usepackage[left=2cm, right=2cm, top=1.5cm, bottom=2.5cm]{geometry}
\usepackage{setspace}

\title{Strassen's algorithm via orbit flip graphs}
\author{Christian Ikenmeyer and Jakob Moosbauer
\\
{\small
University of Warwick}\\
\small{
\texttt{$\{$christian.ikenmeyer, jakob.moosbauer$\}$@warwick.ac.uk}
}}
\date{March 2025}

\newcommand{\ot}{\otimes}
\newcommand{\tensor}{{\textstyle\bigotimes}}
\newcommand{\IF}{\mathbb{F}}
\newcommand{\IN}{\mathbb{N}}

\begin{document}
\raggedbottom

\maketitle

\begin{abstract}
We give a short proof for Strassen's result that the rank of the 2 by 2 matrix multiplication tensor is at most~7. The proof requires no calculations and also no pattern matching or other type of nontrivial verification, and is based solely on properties of a specific order 6 group action. Our proof is based on the recent combination of flip graph algorithms and symmetries.
\end{abstract}

Let $\IF$ be any field, and let $\IF^{2\times 2}$ have the standard basis $\{e_{0,0},e_{0,1},e_{1,0},e_{1,1}\}$.
The $2 \times 2$ matrix multiplication tensor is defined as
$M_2 = \sum_{i,j,k=0}^1 e_{i,j} \ot e_{j,k} \ot e_{k,i} \in \tensor^3(\IF^{2\times 2})$, see \cite[(14.20)]{BCS97}.
The \emph{rank} $R(t)$
for $t \in \tensor^3(\IF^{2\times 2})$
is defined as the smallest number $r\in \IN$ such that there exist $u_{i},v_{i},w_{i}\in \IF^{2\times 2}, i=1,\ldots, r$, with
$t = \sum_{i=1}^{r} u_{i} \ot v_{i} \ot w_{i}$.
Strassen \cite{Str69} proved that $R(M_2)\leq 7$, and we reprove this result.

\begin{spacing}{1.075}
The cyclic group $C_3 = \langle\pi \mid \pi^3=1\rangle$ acts linearly on $\tensor^3(\IF^{2\times 2})$ via $\pi(u\ot v \ot w) = w \ot u \ot v$
and $\IF$-linear extension.
For $i \in \{0,1\}$, let 
$i' := 1-i$.
The cyclic group $C_2= \langle\tau \mid \tau^2=1\rangle$ acts linearly on $\tensor^3(\IF^{2\times 2})$ via
$\tau(e_{i_1,i_2}\ot e_{j_1,j_2} \ot e_{k_1,k_2})
= e_{i'_1,i'_2}\ot e_{j'_1,j'_2} \ot e_{k'_1,k'_2}$
and $\IF$-linear extension.
Both actions commute, so that we have an action of $G := C_3\times C_2$.
We write $G\cdot t := \sum_{g\in G} g(t)$, which is a notation introduced in \cite{BILR19}.
For a standard basis vector $t$ it is easy to visualize
$G\cdot t = G\cdot \tau(t)$ and $G\cdot t = G\cdot \pi(t)$: For $\tau$ we
connect pairs of inverted indices $i$ and $i'$ with red lines, and for $\pi$ we
illustrate the cyclic position rotation with blue lines:
\end{spacing}
\begin{tikzpicture}
\node at (-5,0) {
\begin{minipage}{5cm}
\vspace{-0.3cm}
\begin{align*}
&&G\cdot e_{0,1}\ot e_{1,1} \ot e_{0,0}
\\[0.2cm]
&&=G\cdot e_{1,0}\ot e_{0,0} \ot e_{1,1}
\end{align*}
\end{minipage}
};
\node at (0,-0.1) {and};
\node at (5,0) {
\begin{minipage}{5cm}
\vspace{-0.3cm}
\begin{align*}
&&G\cdot e_{0,1}\ot e_{1,1} \ot e_{0,0}
\\[0.2cm]
&&=G\cdot e_{0,0}\ot e_{0,1} \ot e_{1,1}
\end{align*}
\end{minipage}
};
\draw[thick,red] (-5.27,-0.36) -- (-5.27,0.04);
\draw[thick,red] (-5.04,-0.36) -- (-5.04,0.04);
\draw[thick,red] (-4.29,-0.36) -- (-4.29,0.04);
\draw[thick,red] (-4.06,-0.36) -- (-4.06,0.04);
\draw[thick,red] (-3.32,-0.36) -- (-3.32,0.04);
\draw[thick,red] (-3.09,-0.36) -- (-3.09,0.04);
\draw[thick,blue!50!white] (5.8,-0.30) -- (4.83,0);
\draw[thick,blue!50!white] (4.83,-0.30) -- (6.77,0);
\draw[thick,blue!50!white] (6.77,-0.30) -- (5.8,0);
\end{tikzpicture}

The proof of $R(M_2)\leq 7$ now goes as follows.
There are two different kinds of summands in the 8 summands of~$M_2$: 6 of them form $G\cdot e_{1,0} \ot e_{0,1} \ot e_{1,1}$. For the remaining 2 summands, we can get their sum from a rank~1 tensor with 8 terms and subtracting the 6 superfluous terms as follows:
\[
e_{0,0}\ot e_{0,0} \ot e_{0,0} + e_{1,1}\ot e_{1,1} \ot e_{1,1} \ = \ (e_{0,0}+e_{1,1})\ot (e_{0,0}+e_{1,1})\ot (e_{0,0}+e_{1,1})
- G \cdot e_{0,0}\ot e_{0,0} \ot e_{1,1}.
\]
So far, by naively looking at the summand structure, we have obtained a rank $1+6+6=13$ decomposition for~$M_2$.
We now observe:

\begin{tikzpicture}
\node at (-4.7,1.45) {$M_2 \, - \, (e_{0,0}+e_{1,1})\ot (e_{0,0}+e_{1,1})\ot (e_{0,0}+e_{1,1})$};
\node at (0,0) {\begin{minipage}{\textwidth}
\begin{alignat*}{6}
&=
  G \cdot e_{1,0} \ot e_{0,1} \ot e_{1,1}
 && && 
 &&- G \cdot e_{0,0} \ot e_{0,0} \ot e_{1,1}
\\
&=
 G \cdot e_{1,0} \ot e_{0,1} \ot e_{1,1}
&{} + \big(&- G \cdot e_{0,1} \ot e_{1,1} \ot e_{0,0}&&+\hspace{4.5pt} G \cdot e_{0,1} \ot e_{1,1} \ot e_{0,0} \big)
&&- G \cdot e_{0,0} \ot e_{0,0} \ot e_{1,1}
\\[0.2cm]
&=\mathrlap{
 \big(G \cdot e_{1,0} \ot e_{0,1} \ot e_{1,1}}
&&- G \cdot e_{1,0} \ot e_{0,0} \ot e_{1,1}\big)
&& + \big(G \cdot e_{0,0} \ot e_{0,1} \ot e_{1,1} 
&&- G \cdot e_{0,0} \ot e_{0,0} \ot e_{1,1}\big)
\\
&=
 \mathrlap{\hspace{1.5cm} G \cdot e_{1,0} \ot (e_{0,1}-e_{0,0}) \ot e_{1,1}}
&&
&&+ \mathrlap{\hspace{1.5cm}G \cdot e_{0,0} \ot (e_{0,1}-e_{0,0}) \ot e_{1,1}}
&&
\\
&=
 \mathrlap{\hspace{4.5cm}G \cdot (e_{1,0}+e_{0,0}) \ot (e_{0,1}-e_{0,0}) \ot e_{1,1},}
\end{alignat*}
\end{minipage}};
\draw[thick,red] (-2.27,-0.26) -- (-2.27,0.14);
\draw[thick,red] (-2.04,-0.26) -- (-2.04,0.14);
\draw[thick,red] (-1.29,-0.26) -- (-1.29,0.14);
\draw[thick,red] (-1.06,-0.26) -- (-1.06,0.14);
\draw[thick,red] (-0.32,-0.26) -- (-0.32,0.14);
\draw[thick,red] (-0.09,-0.26) -- (-0.09,0.14);
\draw[thick,blue!50!white] (2.6,-0.20) -- (1.63,0.1);
\draw[thick,blue!50!white] (1.63,-0.20) -- (3.57,0.1);
\draw[thick,blue!50!white] (3.57,-0.20) -- (2.6,0.1);
\end{tikzpicture}
which is of rank at most 6.
Hence, $M_2$ has rank at most~7.

\section*{Context and related work}
For $n\in\IN$, let $M_n := \sum_{i,j,k=0}^{n-1} e_{i,j} \ot e_{j,k} \ot e_{k,i}$.
The matrix multiplication exponent $\omega = \inf\{\log_{n}(R(M_n))\}$ governs the running time of the algorithmic task of multiplying two matrices and of other tasks in linear algebra, for example solving systems of linear equations, see \cite[\S15.1, \S16]{BCS97}.
In order to understand $R(M_n)$ for large~$n$, significant research effort has focused on understanding $R(M_n)$ for small values of $n$, with several papers focusing on $n=2$ \cite{Bre70, Gas71, Fid72, BD73, Pat74, Laf75, Mak75, Yuv78, BC85, Cla88, Cha86, Ale97, GK00, Pat09, GM17, IL19, CILO19}, since no satisfying short proof of $R(M_2)\leq 7$ has been found.
The verification of these proofs usually requires simple but somewhat lengthy calculations, and careful attention to the details has to be paid.
The weaker analogous result for the tensor border rank
also gives $\omega\leq\log_2(7)$ and
can be obtained from Lickteig's analysis of defective secant varieties, see \cite{Lic85}, \cite[\S3.8, \S9]{Lan08}, which is significantly more involved than our proof.

Our new proof of $R(M_2)\leq 7$ is so short that it can hopefully be used
successfully as a blueprint for proofs for other small~$n$.  In fact, it does
arise from a machinery for proofs for small $n$, the flip graph algorithm
\cite{KM23}.  $R(M_n)$ for small values of $n$ has been studied for a long time
\cite{HK71,Lad76,JM86,Mak87,DIS11,Smi13,Sed17,HKS19,HKS21,SS21,Smi21,Fawetal22,Smi23,AIH24,KM25,Sed25}.
The recent flip graph algorithm works as follows: one starts with a simple but
expensive decomposition of $M_n$ and reduces its rank via a series of
simplifications.  Adding ideas from \cite{BILR19} (tensor rank decompositions of
$M_3$), this technique has recently been improved by considering the symmetry
groups $C_3$ and $C_3\times C_2$ to get new upper bounds on $R(M_5)$ and
$R(M_6)$, see \cite{MP25}.  Our proof can be interpreted as a refined version of
the output of this newest flip graph algorithm.  Note that \cite{BILR19} also
consider a group of order 6, but the group there is $C_3 \rtimes C_2 \simeq
S_3$.  The flip graph proof technique requires to specify a trivial but expensive solution,
a group (usually a subgroup of the symmetry group of $M_n$), and a sequence of
trivial simplifications.  As a consequence, our proof generates 7
summands instead of starting with them and verifying their correctness.
\cite{deG78} showed that any rank 7 decomposition is equivalent to Strassen's original decomposition.
It is
an open question which starting solutions and which groups can be used to get
satisfying short proofs for $M_n$, $n \geq 3$.

\bibliographystyle{alpha}
{\footnotesize
\bibliography{lit}
}
\end{document}